How much weight of H2O2 is needed to make 1.5 N of solution in 1L

Grade:Upto college level

1 Answers

Naveen Kumar
9 years ago

n-factor of H2O2(n)=2
Hence, by relation Normality=Molarity*n,
Molarity of H2O2=1.5/2=0.75M
So for 1L, moles of H2O2=0.75M*1L=0.75 mol
i.e. 0.75 mol of H2O2 required for 1L
Molecular weight of H2O2=34g(mol)-1
Hence weight of H2O2 required=0.75*34=25.5g